OpenStudy (anonymous):

solve for t? 1863.3=(1.00675)^4t

OpenStudy (precal):

log both sides

OpenStudy (anonymous):

im still confuse.

OpenStudy (anonymous):

equation please?

OpenStudy (ghazi):

log(1863.3) = log(1.00675^4t)....now log (A^B)= B*log (A)....now you can do it

OpenStudy (precal):

drop the power in front of the log

OpenStudy (ghazi):

yes...it is the property of log

OpenStudy (anonymous):

thankss

OpenStudy (precal):

hope that helps, if not let me know